# Intrinsic training dynamics of deep neural networks

### v1

@@ 3.3 Application to general deep ReLU networks
This leads to the following corollary (proved in Appendix˜F) which guarantees the existence of a maximal set of conservation laws big enough to ensure the intrinsic recoverability property.
@@
There exists a dense open set $\Theta$ of $\mathbb{R}^{D}$ such that any $\theta_{0}\in\Theta$ admits an open neighborhood $U\subseteq\Theta$ on which $\theta_{0}$ satisfies the intrinsic recoverability property, and thus the intrinsic dynamic property with respect to $\phi_{\mathtt{ReLU}}$.
In practice, the known conservation laws given in Proposition˜2.9 yield, on a dense open subset, $m$ independent conservation laws, where $m$ corresponds to the number of hidden neurons. To verify that these are indeed the only ones, one must check that the trace of $\mathbb{W}_{\phi_{\mathtt{ReLU}}}$ has dimension DmD-m; while we do not prove this here, it is empirically supported by Marcotte et al. (2023), which confirms that $\mathrm{Lie}\big{(}\mathbb{W}_{\phi_{\mathtt{ReLU}}}\big{)}(\theta)=\mathbb{W}_{\phi_{\mathtt{ReLU}}}(\theta)$ has dimension DmD-m when sampling random values of $\theta$, as well as random dimensions and depths.
As a concrete example the following proposition (proved in Appendix˜G) provides the first closed form formula of the intrinsic dynamic for a three-layer ReLU network with scalar input and output.
@@
with $u\in\mathbb{R}^{n}$, $V\in\mathbb{R}^{n\times m}$, and $w\in\mathbb{R}^{m}$. Define $\Theta:=\{(u,V,w):u_{i},V_{ij},w_{j}\neq 0\ \forall i,j\}$, and let $\theta(t)$ be the maximal solution to (1) with $\theta(0)=\theta_{0}\in\Theta$. The flow preserves the $n+m$ conservation laws $\mathbf{h}(\theta):=\big{(}(u_{i}^{2}-\sum_{j}V_{ij}^{2})_{i=1}^{n},(w_{j}^{2}-\sum_{i}V_{ij}^{2})_{j=1}^{m}\big{)}$, and the intrinsic dynamics $\dot{z}=-K_{\theta_{0}}(z)\nabla f(z)$ on $z=\mathtt{vec}(Z)$ corresponds to
$\dot{Z}=-\mathrm{ddiag}(\nabla f(Z)Z^{\top})\,\mathrm{diag}(\boldsymbol{\alpha})^{-1}\,Z-\mathrm{diag}(\boldsymbol{\alpha})\,\nabla f(Z)\,\mathrm{diag}(\boldsymbol{\beta})-Z\,\mathrm{diag}(\boldsymbol{\beta})^{-1}\,\mathrm{ddiag}(Z^{\top}\nabla f(Z)),$
where: a) for any matrix $M$, we denote $\mathrm{ddiag}(M)=\mathrm{diag}\bigl{(}\mathrm{Diag}(M)\bigr{)}$, where $\mathrm{Diag}(M)$ extracts its diagonal as a vector and $\mathrm{diag}(v)$ denotes the diagonal matrix with entries of $v$; and b) the vectors $\boldsymbol{\alpha}=\boldsymbol{\alpha}(Z,\mathbf{h}(\theta_{0}))\in\mathbb{R}_{>0}^{n}$ and $\boldsymbol{\beta}:=\boldsymbol{\beta}(Z,\mathbf{h}(\theta_{0}))\in\mathbb{R}_{>0}^{m}$ (uniquely determined by $Z$ and $\mathbf{h}(\theta_{0})$) satisfy
$\boldsymbol{\alpha}^{2}-|Z|^{2}\,\mathrm{diag}(\boldsymbol{\beta})^{-1}\mathbf{1}_{n}-\boldsymbol{\lambda}\odot\boldsymbol{\alpha}=0,\quad\boldsymbol{\beta}^{2}-(|Z|^{2})^{\top}\,\mathrm{diag}(\boldsymbol{\alpha})^{-1}\mathbf{1}_{m}-\boldsymbol{\mu}\odot\boldsymbol{\beta}=0,$ (10)
with $|Z|^{2}\in\mathbb{R}^{n\times m}$ the element-wise square on the matrix $Z\in\mathbb{R}^{n\times m}$) and with $\boldsymbol{\lambda}\in\mathbb{R}^{n}$, $\boldsymbol{\mu}\in\mathbb{R}^{m}$ such that $\mathbf{h}(\theta_{0})=(\boldsymbol{\lambda},\boldsymbol{\mu})$. When $\boldsymbol{\lambda},\boldsymbol{\mu}=0$, (10) entirely characterizes $(\boldsymbol{\alpha},\boldsymbol{\beta})$.
4 Deep linear neural networks and linear neural ODEs
For $L$-layer linear networks, $\theta=(U_{1},\ldots,U_{L})$ and the path-lifting formalism (Gonon et al., 2023) yields a factorization via $\phi_{\mathtt{ReLU}}$, leading to an intrinsic dynamics by the results of the previous section. It is more common however to consider the dynamics of $Z_{L}:=\phi_{\mathrm{Lin}}(\theta_{L})=U_{L}\cdots U_{1},$ since $\phi_{\mathtt{lin}}$ is more efficient that $\phi_{\mathtt{ReLU}}$ in terms of dimension reduction. We now analyze the dynamics of $Z_{L}(t)$. The gradient flow $\dot{\theta}_{L}=-\nabla\ell(\theta_{L})$ gives the evolution of $Z_{L}$ (see e.g. (Bah et al., 2022, Lemma 2)):
@@ 4 Deep linear neural networks and linear neural ODEs
Observe that the metric $M(\theta_{L})$ on $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}z_{L}=\mathtt{vec}(Z_{L})$} is entirely characterized by the functions
$(S_{j}(\theta_{L}),T_{j+1}(\theta_{L}))_{j=1}^{L-1}.$ (12)
Definition 4.1 (Relaxed balanced conditions).
We say that $\theta_{L}\coloneqq(U_{L},\cdots,U_{1})$ satisfies the relaxed balanced condition if there exists $\boldsymbol{\lambda}\coloneqq(\lambda_{i})_{i}\in\mathbb{R}^{L-1}$ such that
@@
(0-)balanced conditions (Bah et al., 2022, Def 1) (Arora et al., 2019, Def 1) correspond to $\boldsymbol{\lambda}=0$.
@@
It is worth noting that Dominé et al. (2025) used this exact same condition and called it the $\lambda$-balanced condition. However, the definition of $\lambda$-balanced condition is already used (see (Arora et al., 2019, Def 1)) by the literature to refer to the weaker condition $\|U_{i+1}^{\top}U_{i+1}-U_{i}U_{i}^{\top}\|\leq{\lambda_{i}}.$ Other works (see e.g. Tarmoun et al. (2021); Braun et al. (2022); Varre et al. (2023)) use stronger conditions on the initializations, that satisfy in particular the relaxed balanced conditions of Definition˜4.1.
4.1 Deep linear neural networks
@@
Matrix factorization. We consider the two-layer case where $\theta\coloneqq(U,V)\in\mathbb{R}^{n\times r}\times\mathbb{R}^{m\times r}$ and with $Z=\phi_{\mathtt{Lin}}(\theta)\coloneqq UV^{\top}\in\mathbb{R}^{n\times m}$. We assume $\theta(t)$ satisfies the gradient flow (1) with $\theta(0)=(U_{t=0},V_{t=0})$. We denote $S\coloneqq U_{t=0}^{\top}U_{t=0}-V_{t=0}^{\top}V_{t=0}\in\mathbb{R}^{r\times r}$.
If $\theta_{0}$ satisfies the balanced condition (13) $S=0$, then (Arora et al., 2018, Theorem 1) (Bah et al., 2022, Lemma 2) $\theta_{0}$ satisfies the intrinsic metric property with respect to $\phi_{\mathtt{Lin}}$ and
@@
We now generalize this result (see Appendix˜H for a proof) to a broader class of initializations: all initializations satisfying the relaxed balanced condition (13) possess the intrinsic metric property.
@@
$\dot{Z}=-\Pi_{ZZ^{\top}}\left[\frac{\lambda}{2}\mathrm{Id}_{n}+\frac{1}{2}\sqrt{\lambda^{2}\mathrm{Id}_{n}+4\,ZZ^{\top}}\right]\nabla f(X)-\nabla f(X)\Pi_{Z^{\top}Z}\left[-\frac{\lambda}{2}\mathrm{Id}_{m}+\frac{1}{2}\sqrt{\lambda^{2}\mathrm{Id}_{m}+4\,Z^{\top}Z}\right],$ (15)
@@
Note that (14) corresponds indeed to (15) with $\lambda=0$. Note also that Theorem˜4.3 generalizes to the case $r\leq\text{min}(n,m)$ the expression obtained in (Dominé et al., 2025, Theorem 5.2) for the special case $r=\text{min}(n,m)$ (if Dominé et al. (2025) focus in general on the squared loss, the proof of their Theorem 5.2 does not rely on the use of this specific loss: this result can be applied for any loss, as ours).
The following theorem shows that the relaxed balanced condition is actually a necessary condition when $r\leq\text{max}(n,m)$ to have the intrinsic metric property. Its proof (see Appendix˜I) relies on showing the non-inclusion of the kernels of (7).
@@
Let $\theta_{0}\coloneqq(U_{t=0},V_{t=0})$. Assume that both $U_{t=0}\in\mathbb{R}^{n\times r}$ and $V_{t=0}\in\mathbb{R}^{m\times r}$ have a full rank and that $r\leq\text{max}(n,m)$. If $S\coloneqq U_{t=0}^{\top}U_{t=0}-V_{t=0}^{\top}V_{t=0}\neq\lambda\mathrm{Id}_{r}$, then $\theta_{0}$ does not satisfy the intrinsic metric property (Definition˜2.10) with respect to $\phi_{\mathtt{Lin}}$.
The case when $r>\text{max}(n,m)$ is still an open question. In the particular case $n=m=1$, for any $r$, the following proposition (see Appendix˜J for a proof) shows that all initializations do satisfy the intrinsic metric property with respect to $\phi_{\mathtt{Lin}}$.
@@
Deep linear neural networks. We now consider linear neural networks of arbitrary depth, with square weight matrices $\theta_{L}\coloneqq(U_{L},\dots,U_{1})$ where $U_{i}\in\mathbb{R}^{n\times n}$. The following theorem (see Appendix˜K for a proof) generalizes Theorem˜4.3 to this setting.
@@
If $\theta_{L}(0)$ satisfies the relaxed balanced condition (Definition˜4.1) with $\boldsymbol{\lambda}=(\lambda_{i})_{i}$ then during the trajectory $\theta_{L}(t)$ of (1), the matrices in (11) satisfy $S_{j}(\theta_{L}(t))=Q_{j}(U_{L}(t)U_{L}(t)^{\top})$ and $T_{j}(\theta_{L}(t))=R_{j}(U_{1}(t)^{\top}U_{1}(t))$, where $Q_{j}(x):=\prod_{k=0}^{L-j-1}(x-a_{k})$ with $a_{0}:=0$ and $a_{k}:=\sum_{i=1}^{k}\lambda_{L-i}$ for $k=1,\cdots L-1$ and $R_{j}(x):=\prod_{k=0}^{j-2}(x-b_{k})$ with $b_{0}:=0$ and $b_{k}:=-\sum_{i=1}^{k}\lambda_{i}$. Moreover $U_{L}U_{L}^{\top}$ (resp. $U_{1}^{\top}U_{1}$) is the unique root of $Z_{L}Z_{L}^{\top}=Q_{0}(U_{L}U_{L}^{\top})$ (resp. of $Z_{L}^{\top}Z_{L}=R_{L-1}(U_{1}^{\top}U_{1})$) with spectrum lower bounded by $\max_{0\leq k\leq L-1}a_{k}$ (resp. by $\max_{0\leq k\leq L-2}b_{k}$). This implies that all matrices in (12) are entirely characterized by $Z_{L}$ and the initialization, hence $\theta_{L}(0)$ satisfies the intrinsic dynamic property on $\mathbb{R}^{D}$ with respect to $\phi_{\mathtt{Lin}}$.
In the particular case of balanced conditions ($\boldsymbol{\lambda}=0$), our theorem recovers the dynamics described in (Arora et al., 2018, Theorem 1), (Bah et al., 2022, Lemma 2).
Network type Parametrization $\boldsymbol{\phi_{\text{Lin}}}$ $\boldsymbol{\phi_{\text{ReLU}}}$
Linear network IMP: only for relaxed balanced $\theta_{0}$ Dimension: $d^{\prime}\boldsymbol{\ll}D$ IMP: for all $\theta_{0}$ $d^{\prime}=D-k\boldsymbol{\approx}D$
DAG ReLU N/A IMP: for all $\theta_{0}$ $d^{\prime}=D-k\boldsymbol{\approx}D$
4.2 Infinitely deep linear networks

### v2

@@ 3 Application for general ReLU networks and Linear Networks
The following theorem (proved in Section˜F.2 for general DAG ReLU networks) refines this result by allowing some zero entries in the parameters.
Theorem 3.2.
Consider an MLP with at least two neurons in each layer (including input and output layers), and denote $\Theta\subseteq\mathbb{R}^{D}$ the set of all parameters with at most one zero component. On this set, which is open, dense, and with a single connected component, the parameterization $\phi_{\mathtt{ReLU}}$ of Gonon et al. (2024) satisfies the Frobenius property.
This leads to the following corollary (proved under weaker assumptions in Section˜F.3) which guarantees the existence of a maximal set of conservation laws big enough to ensure the intrinsic recoverability property.
@@
For a general DAG ReLU network (resp. for an MLP with at least two neurons per layer), every element $\theta$ of the set $(\mathbb{R}\setminus\{0\})^{D}$ (resp. of the set $\Theta$ of Theorem˜3.2) admits an open neighborhood $U$ in this set on which the intrinsic recoverability property (and thus the intrinsic dynamic property with respect to $\phi_{\mathtt{ReLU}}$) holds.
Moreover, the known conservation laws given in Proposition˜2.9 are independent on $\Theta$ from Theorem˜3.2 (actually on an even larger set containing $\Theta$). Thus, they yield $m$ independent conservation laws on $\Theta$, where $m$ corresponds to the number of hidden neurons. To verify that these are indeed the only ones, one must check that the trace of $\mathbb{W}_{\phi_{\mathtt{ReLU}}}$ has dimension DmD-m; while it is empirically supported by Marcotte et al. (2023), which confirms that $\mathrm{Lie}\big(\mathbb{W}_{\phi_{\mathtt{ReLU}}}\big)(\theta)=\mathbb{W}_{\phi_{\mathtt{ReLU}}}(\theta)$ has dimension DmD-m when sampling random values of $\theta$, as well as random dimensions and depths, the following corollary establishes it theoretically (see Section˜F.4 for a proof).
Corollary 3.4.
Let us consider a MLP architecture with at least two neurons per layer, and with $m$ hidden neurons. Then the $m$ independent conservation laws for $\phi_{\texttt{ReLU}}$ given in Proposition˜2.9 are exhaustive on the set $\Theta$ from Theorem˜3.2: there is no more conservation laws than these ones.
As a concrete example the following proposition (proved in Appendix˜G) provides the first closed-form formula for the intrinsic dynamic for a three-layer ReLU network with scalar input and output.
@@
with $u\in\mathbb{R}^{n}$, $V\in\mathbb{R}^{n\times m}$, and $w\in\mathbb{R}^{m}$. Define $\Theta:=\{(u,V,w):u_{i},V_{ij},w_{j}\neq 0\ \forall i,j\}$, and consider the $n+m$ conservation laws $\mathbf{h}(\theta):=\big((u_{i}^{2}-\sum_{j}V_{ij}^{2})_{i=1}^{n},(w_{j}^{2}-\sum_{i}V_{ij}^{2})_{j=1}^{m}\big)$ for $\phi_{\texttt{ReLU}}$. Every $\theta_{0}\in\Theta$ satisfies the intrinsic dynamics with respect to $\phi_{\texttt{ReLU}}$, which reads $\dot{z}=-K_{\theta_{0}}(z)\nabla f(z)$ with $z=\mathtt{vec}(Z)$ corresponding to
$\dot{Z}=-\mathrm{ddiag}(\nabla f(Z)Z^{\top})\,\mathrm{diag}(\bm{\alpha})^{-1}\,Z-\mathrm{diag}(\bm{\alpha})\,\nabla f(Z)\,\mathrm{diag}(\bm{\beta})-Z\,\mathrm{diag}(\bm{\beta})^{-1}\,\mathrm{ddiag}(Z^{\top}\nabla f(Z)),$
where: a) for any matrix $M$, $\mathrm{ddiag}(M):=\mathrm{diag}\bigl(\mathrm{Diag}(M)\bigr)$, where $\mathrm{Diag}(M)$ extracts its diagonal as a vector and $\mathrm{diag}(v)$ is the diagonal matrix with entries of $v$; and b) the vectors $\bm{\alpha}=\bm{\alpha}(Z,\mathbf{h}(\theta_{0}))\in\mathbb{R}_{>0}^{n}$ and $\bm{\beta}:=\bm{\beta}(Z,\mathbf{h}(\theta_{0}))\in\mathbb{R}_{>0}^{m}$ (uniquely determined by $Z$ and $\mathbf{h}(\theta_{0})$) satisfy
$\bm{\alpha}^{2}-|Z|^{2}\,\mathrm{diag}(\bm{\beta})^{-1}\mathbf{1}_{n}-\bm{\lambda}\odot\bm{\alpha}=0,\quad\bm{\beta}^{2}-(|Z|^{2})^{\top}\,\mathrm{diag}(\bm{\alpha})^{-1}\mathbf{1}_{m}-\bm{\mu}\odot\bm{\beta}=0,$ (10)
with $|Z|^{2}\in\mathbb{R}^{n\times m}$ the element-wise square on the matrix $Z\in\mathbb{R}^{n\times m}$ and with $\bm{\lambda}\in\mathbb{R}^{n}$, $\bm{\mu}\in\mathbb{R}^{m}$ such that $\mathbf{h}(\theta_{0})=(\bm{\lambda},\bm{\mu})$. When $\bm{\lambda},\bm{\mu}=0$, equation 10 entirely characterizes $(\bm{\alpha},\bm{\beta})$.
3.1 Deep linear neural networks and linear neural ODEs
For $L$-layer linear networks, $\theta=(U_{1},\ldots,U_{L})$ and the path-lifting formalism (Gonon et al., 2024) yields a factorization via $\phi_{\mathtt{ReLU}}$, leading to an intrinsic dynamic by the results of the previous section. It is more common, however, to consider the dynamics of $Z_{L}:=\phi_{\mathrm{Lin}}(\theta_{L})=U_{L}\cdots U_{1},$ since $\phi_{\mathtt{Lin}}$ is more efficient than $\phi_{\mathtt{ReLU}}$ in terms of dimension reduction. We now analyze the dynamics of $Z_{L}(t)$. The gradient flow $\dot{\theta}_{L}=-\nabla\ell(\theta_{L})$ gives the evolution of $Z_{L}$ (see e.g. (Bah et al., 2022, Lemma 2)):
@@ 3.1 Deep linear neural networks and linear neural ODEs
The metric $M(\theta_{L})$ on $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}z_{L}=\mathtt{vec}(Z_{L})$} is thus entirely characterized by $(S_{j}(\theta_{L}),T_{j+1}(\theta_{L}))_{j=1}^{L-1}$.
Definition 3.6 (Relaxed balanced condition).
We say that $\theta_{L}\coloneqq(U_{L},\cdots,U_{1})$ satisfies the relaxed balanced condition if there exists $\bm{\lambda}\coloneqq(\lambda_{i})_{i}\in\mathbb{R}^{L-1}$ such that
@@
(0-)balanced conditions (Bah et al., 2022, Def 1) (Arora et al., 2019, Def 1) correspond to $\bm{\lambda}=0$.
@@
It is worth noting that Dominé et al. (2025) used this exact same condition and called it the $\lambda$-balanced condition. However, the definition of $\lambda$-balanced condition is already used (see (Arora et al., 2019, Def 1)) by the literature to refer to the weaker condition $\|U_{i+1}^{\top}U_{i+1}-U_{i}U_{i}^{\top}\|\leq{\lambda_{i}}.$ Other works (see e.g. Tarmoun et al. (2021); Braun et al. (2022); Varre et al. (2023)) use stronger conditions on the initializations, that in particular satisfy the relaxed balanced conditions of Definition˜3.6.
3.1.1 Deep linear neural networks
@@
Matrix factorization. We consider the two-layer case where $\theta\coloneqq(U,V)\in\mathbb{R}^{n\times r}\times\mathbb{R}^{m\times r}$ and with $Z=\phi_{\mathtt{Lin}}(\theta)\coloneqq UV^{\top}\in\mathbb{R}^{n\times m}$. We assume $\theta(t)$ satisfies the gradient flow equation 1 with $\theta(0)=(U_{t=0},V_{t=0})$. We denote $S\coloneqq U_{t=0}^{\top}U_{t=0}-V_{t=0}^{\top}V_{t=0}\in\mathbb{R}^{r\times r}$. If $\theta_{0}$ satisfies the balanced condition equation 12 $S=0$, then (Arora et al., 2018, Theorem 1) (Bah et al., 2022, Lemma 2) $\theta_{0}$ satisfies the intrinsic metric property with respect to $\phi_{\mathtt{Lin}}$ and
@@
We generalize this result (see Appendix˜H for a proof) to a broader class of initializations: all initializations satisfying the relaxed balanced condition equation 12 possess the intrinsic metric property.
@@
$\dot{Z}=-\Pi_{Z}\left[\frac{\lambda}{2}\mathrm{Id}_{n}+\frac{1}{2}\sqrt{\lambda^{2}\mathrm{Id}_{n}+4\,ZZ^{\top}}\right]\nabla f(Z)-\nabla f(Z)\Pi_{Z^{\top}Z}\left[-\frac{\lambda}{2}\mathrm{Id}_{m}+\frac{1}{2}\sqrt{\lambda^{2}\mathrm{Id}_{m}+4\,Z^{\top}Z}\right],$ (14)
@@
Note that equation 13 corresponds indeed to equation 14 with $\lambda=0$ and that $r\leq\min(n,m)$ is necessary to have $S=\lambda\mathrm{Id}_{r}$ for some $\lambda\neq 0$. Note also that Theorem˜3.8 generalizes to the case $r\leq\text{min}(n,m)$ the expression obtained in (Dominé et al., 2025, Theorem 5.2) for the special case $r=\text{min}(n,m)$ (if Dominé et al. (2025) focus in general on the squared loss, the proof of their Theorem 5.2 does not rely on the use of this specific loss: this result can be applied for any loss, as ours). The following theorem shows that the relaxed balanced condition is actually a necessary condition when $r\leq\text{max}(n,m)$ to have the strong intrinsic dynamic property (see Definition˜2.10 in Appendix˜I), a variant of Definition˜2.6 where $f$ is piecewise $\mathcal{C}^{2}$. Its proof (see Appendix˜I) relies on showing the non-inclusion of the kernels of equation (ii).
@@
Let $\theta_{0}\coloneqq(U_{t=0},V_{t=0})$. Assume that both $U_{t=0}\in\mathbb{R}^{n\times r}$ and $V_{t=0}\in\mathbb{R}^{m\times r}$ have full rank and that $r\leq\text{max}(n,m)$. If $S\coloneqq U_{t=0}^{\top}U_{t=0}-V_{t=0}^{\top}V_{t=0}\neq\lambda\mathrm{Id}_{r}$, then $\theta_{0}$ does not satisfy the strong intrinsic dynamic property with respect to $\phi_{\mathtt{Lin}}$.
The case $r>\text{max}(n,m)$ is still open. For $n=m=1$ and any $r$, the following proposition (proved in Appendix˜J) shows that all initializations do satisfy the intrinsic metric property with respect to $\phi_{\mathtt{Lin}}$.
@@
Deep linear neural networks. Consider linear neural networks of arbitrary depth, with square weight matrices $\theta_{L}\coloneqq(U_{L},\dots,U_{1})$, $U_{i}\in\mathbb{R}^{n\times n}$. Indeed, this is precisely the main case where a non-trivial relaxed balanced condition ($\lambda_{i}\neq 0$) can occur at every layer $i$, since $n_{i}=n$ and thus $n_{i}\leq\min(n_{i-1},\,n_{i+1})$ for all $i$. The following theorem (proved in Appendix˜K) generalizes Theorem˜3.8 to this setting. In the case of balanced conditions ($\bm{\lambda}=0$), our theorem recovers the dynamics described in (Arora et al., 2018, Theorem 1), (Bah et al., 2022, Lemma 2).
@@
If $\theta_{L}(0)$ satisfies the relaxed balanced condition (Definition˜3.6) with $\bm{\lambda}=(\lambda_{i})_{i}$ then during the trajectory $\theta_{L}(t)$ of equation 1, the matrices in equation 11 satisfy $S_{j}(\theta_{L}(t))=Q_{j}(U_{L}(t)U_{L}(t)^{\top})$ and $T_{j}(\theta_{L}(t))=R_{j}(U_{1}(t)^{\top}U_{1}(t))$, where $Q_{j}(x):=\prod_{k=0}^{L-j-1}(x-a_{k})$ with $a_{0}:=0$ and $a_{k}:=\sum_{i=1}^{k}\lambda_{L-i}$ for $k=1,\cdots L-1$ and $R_{j}(x):=\prod_{k=0}^{j-2}(x-b_{k})$ with $b_{0}:=0$ and $b_{k}:=-\sum_{i=1}^{k}\lambda_{i}$. Moreover $U_{L}U_{L}^{\top}$ (resp. $U_{1}^{\top}U_{1}$) is the unique root of $Z_{L}Z_{L}^{\top}=Q_{0}(U_{L}U_{L}^{\top})$ (resp. of $Z_{L}^{\top}Z_{L}=R_{L-1}(U_{1}^{\top}U_{1})$) with spectrum lower bounded by $\max_{0\leq k\leq L-1}a_{k}$ (resp. by $\max_{0\leq k\leq L-2}b_{k}$). This implies that all matrices in equation 11 are entirely characterized by $Z_{L}$ and the initialization, hence $\theta_{L}(0)$ satisfies the intrinsic dynamic property on $\mathbb{R}^{D}$ with respect to $\phi_{\mathtt{Lin}}$.
3.1.2 Infinitely deep linear networks